# LLM-Powered Agents for Navigating Venice’s Historical Cadastre

### v1

@@ 14 How many unique property types are represented throughout the dataset?
15 What is the total count of "casa" properties listed?
16 What is the total number of "casa" present in the dataset?
17 What is the count of properties designated as "casa"?
18 What is the count of dataset properties that are classified as types other than "altro appartamento"?
19 How many "casa in soler" properties exist in total?
@@ 20 Which property generates the least amount of rent income?
21 What is the total rent revenue generated from properties of the "bottega da casarol" variety?
@@ 22 Who are the proprietors of estates with a rent revenue of 30?
23 What is the aggregate number of properties that the "rissardi" family members possess?
@@ 24 What is the average number of properties for each property type?
25 How much rent do properties on "calle delle carozze" generate altogether?
26 What is the total income from rent for all properties classified as "casa"?
@@ 28 What is the count of properties that are being leased for an amount below the average rent income?
29 What are the various types of properties located on "calle della torre"?
@@ 30 What is the number of properties yielding a rental income lesser than the dataset’s average?
31 How much rental income do properties typically generate in "al ponte di san provolo" on average?
32 How much rent do all the properties combined on "fondamenta san domenico" generate?
33 What variety of property can be found in "teren alli gesouiti"?
34 What are the names of the families who possess land in "rio san baseggio"?
35 Who holds ownership of properties classified as "magazen due"?
@@ 39 Who are the proprietors possessing real estate with rents fluctuating between 20 to 100 ducati?
40 Can you tell me about the spectrum of rents collected on "calle de franchi"?
@@ 41 What is the average rental income generated by each family?
42 How much rental income is typically received from properties located outside of "segue rughetta verso calle del fontico" on average?
43 How much income does "carlo" "gritti" generate from rent on his property?
@@ 44 Who is the top earner from a single property in terms of rent and what is the amount?
45 Who holds ownership of the estates situated at "corte de cà celsi"?
46 What variety of holdings does "marin" "bernardo" possess?
47 Could you enumerate the various types of real estate that "marin" "bernardo" possesses?
@@ 48 Which kind of property is associated with the lowest average rental earnings?
49 Can you list the various property categories under "francesco" "patarol"’s ownership?
50 What is the count of properties located on "la calle vicina al campiel dal panizza in arzere"?
@@ 51 How many categories of properties are solely found in certain locations?
52 What is the sum of rental revenue that "francesco" "mosto" collects from his entire real estate portfolio?
53 What types of properties can be found in "calle di santa cattarina principia alle fondamenta nove"?
@@ 56 Which properties generate a rental income exceeding 50 ducati?
57 In each location, what is the count of "casa" properties?
@@ 65 Who are the individuals possessing properties in several areas?
66 Is there availability of "casa" type properties located on "calle della malvasia"?
67 Who are the proprietors of a "bottega" located at "fondamenta in fazza la beccaria"?
68 How many properties out of a hundred are found in "fondamenta de carmini"?
69 What’s the minimum amount "zuane" "panizza" earns from the property at "traghetto di san silvestro"?
70 What’s the typical rental income for each type of property owned by "perina" "capello"?
71 What is the count of properties classified under "casa a pepian", and how do they contribute to the overall rental revenue percentage-wise in the data?
@@ 72 In the dataset, which entity holds ownership over the largest number of properties?
73 Does "antonio" possess any real estate on "calle dei ragusei"?
@@ 75 How is the dataset broken down by property type?
76 What is the count of proprietors possessing a minimum of one "porzione di bottega" property?
@@ 78 Who is the principal owner of the most varied property categories?
79 Is a property in "campiello della fraterna" owned by "iseppo maria" "gallo"?
80 Who holds the title as the largest landholder in "sotto le collonelle"?
81 How many properties do owners have on average in the location of "corte carli"?
@@ 82 What’s the total of property owners with investments spread across several areas?
83 In "loco vicino la calle del paradiso", are there identical property types under the ownership of different families?
84 Is the property yielding the least rental income in the dataset owned by "domenico" "michiel"?
85 How does the total earnings from rent of the "casetta, vuota, di solito affittata" stand in comparison with that from different kinds of properties?
@@ 87 How much total rent income do properties with more than 10 ducati generate?
88 How many families own property in "calle corrente dell’occa"?
89 What is the lowest income of "francesco" "giustinian"?
90 For rental properties, what share of the income is derived from those labeled "bottega da fabro"?
@@ 96 How many owners have more than one type of property?
97 How is rent income distributed among properties in "rio terrà"?
@@ A.2 SQL CodeS-7B Prompts
A.3 Text-to-Python Implementation Details
@@
A.4 Text-to-Python Agent prompts
This section contains all the prompts used at each step of the Text-to-Python agent’s process. The first two prompts serve as system prompts, provided to the agent along with a description of the datasets as part of its context. The subsequent prompts align with the agent’s workflow, as shown in Figure 5. The process begins with extracting references, followed by entity extraction. Next, the agent generates a plan and subsequently writes the necessary code. If errors occur during code execution, the agent debugs the code based on the Python console’s error messages. In the following prompts, all inputs used to generate prompts are highlighted in blue. In context examples are given for reference and entity extraction.
@@ A.4 Text-to-Python Agent prompts
Appendix B Additional Results
B.1 Closed vs Open Source model
Figure 8: Performance comparison between GPT4-o and Llama3-70b
B.2 Qualitative analysis of data operations
Table B.2 illustrates how a language model agent systematically converts complex questions in natural language into actionable data analysis operations. Each question is deconstructed into key entities and references, mapping natural language elements (e.g., “rent price,” “square”) to corresponding dataset columns and types. Subsequently, the agent determines the appropriate analytical procedure, such as correlation analysis, filtering, or aggregation tailored to the query. For example, a question about the correlation between rent price and proximity to squares in 1740 is translated into a Pearson correlation calculation. Similarly, queries involving temporal comparisons or categorical relationships leverage statistical methods like chi-square tests or comparative metrics
Questions Entity & Reference Code
Is the building rent price correlated with the distance from the nearest square in 1740? “square“ $\rightarrow$ type in Landmarks “rent price“ $\rightarrow$ price in Catastici_1740 Pearson correlation between rent price and distance to nearest square
Which neighborhood contains the buildings with the highest rent price on average in 1740? “rent price“ $\rightarrow$ price in Catastici_1740 Uses “parish“ as “neighborhood“ and finds the parish with the highest average rent price.
How many people live within 100 meters from the closest church in 1808? ‘Church“ $\rightarrow$ type in Landmarks “people“ $\rightarrow$ owner name in Sommarionni_1808 Filters buildings within 100 meters of a church and counts unique people.
On average, are buildings with multiple functions more expensive than the ones with a single function in 1740? “single function“, “multiple function“ $\rightarrow$ function “more expensive“ $\rightarrow$ price in Catastici_1740 Separates single and multiple function buildings, compares average rent prices.
Which square is surrounded with the buildings with the most diverse functions within 100 meters in 1740? “square“ $\rightarrow$ type in Landmarks “diverse function“ $\rightarrow$ function in Catastici_1740 Counts buildings within 100 meters of squares, outputs the square with the most diverse building functions.
What is the average distance between workshops and their nearest square in 1740? “workshop“ $\rightarrow$ [’magazzeno’, ’orto’] in function of Catastici_1740 Filters workshops and calculates their average distance to the nearest square.
Is there any correlation between the family name and the profession in 1740? “family name“ and “profession“ $\rightarrow$ profession in Catastici_1740 Creates contingency table and uses chi-square test of independence to check for correlation.
In which parish do lawyers own the most number of buildings in 1740? “parish“ $\rightarrow$ parish in Catastici_1740 and “laywers“ $\rightarrow$ [’avocato’, ’avvocato’] in profession column Uses "parish" as "district", counts buildings per parish, outputs the family with most buildings on average.
How many medical doctors are there in Venice in 1740? “medical doctor“‘ $\rightarrow$ medico in profession of Catastici_1740 Filters dataset for "medico" profession and counts the number of medical doctors.
Did the number of buildings with more than one function increase over time from 1740 to 1808? “more than one function“ $\rightarrow$ function count in Catastici_1740 and Sommarioni_1808 Compares the number of multi-function buildings in 1740 and 1808.
Which family increased the number of buildings they owned the most from 1740 to 1808? “family“ $\rightarrow$ name in Catastici_1740 and Sommarioni_1808 Counts the number of buildings per family in both datasets, outputs the family with the highest increase.
How many new families appeared in Venice in 1808 that were not present in 1740? “new families“ $\rightarrow$ name in Catastici_1740 and Sommarioni_1808 Counts the number of families present in 1808 but not in 1740.
Table 4: Analysis of consistent answers. The entity extraction is expressed using the following grammar:“element of question“ is mapped to ($\rightarrow$) a column in a Dataset
B.3 Execution Traces
B.3.1 Successful Trace
Figure 9: Example of a successful executation trace of the text-to-Python agent
B.3.2 Successful Trace
Figure 10: Example of a failing executation trace of the text-to-Python agent

### v2

@@ 14 How many unique property types are represented throughout the dataset?
15 What is the total count of ”casa” properties listed?
16 What is the total number of ”casa” present in the dataset?
17 What is the count of properties designated as ”casa”?
18 What is the count of dataset properties that are classified as types other than ”altro appartamento”?
19 How many ”casa in soler” properties exist in total?
@@ 20 Which property generates the least amount of rent income?
21 What is the total rent revenue generated from properties of the ”bottega da casarol” variety?
@@ 22 Who are the proprietors of estates with a rent revenue of 30?
23 What is the aggregate number of properties that the ”rissardi” family members possess?
@@ 24 What is the average number of properties for each property type?
25 How much rent do properties on ”calle delle carozze” generate altogether?
26 What is the total income from rent for all properties classified as ”casa”?
@@ 28 What is the count of properties that are being leased for an amount below the average rent income?
29 What are the various types of properties located on ”calle della torre”?
@@ 30 What is the number of properties yielding a rental income lesser than the dataset’s average?
31 How much rental income do properties typically generate in ”al ponte di san provolo” on average?
32 How much rent do all the properties combined on ”fondamenta san domenico” generate?
33 What variety of property can be found in ”teren alli gesouiti”?
34 What are the names of the families who possess land in ”rio san baseggio”?
35 Who holds ownership of properties classified as ”magazen due”?
@@ 39 Who are the proprietors possessing real estate with rents fluctuating between 20 to 100 ducati?
40 Can you tell me about the spectrum of rents collected on ”calle de franchi”?
@@ 41 What is the average rental income generated by each family?
42 How much rental income is typically received from properties located outside of ”segue rughetta verso calle del fontico” on average?
43 How much income does ”carlo” ”gritti” generate from rent on his property?
@@ 44 Who is the top earner from a single property in terms of rent and what is the amount?
45 Who holds ownership of the estates situated at ”corte de cà celsi”?
46 What variety of holdings does ”marin” ”bernardo” possess?
47 Could you enumerate the various types of real estate that ”marin” ”bernardo” possesses?
@@ 48 Which kind of property is associated with the lowest average rental earnings?
49 Can you list the various property categories under ”francesco” ”patarol”’s ownership?
50 What is the count of properties located on ”la calle vicina al campiel dal panizza in arzere”?
@@ 51 How many categories of properties are solely found in certain locations?
52 What is the sum of rental revenue that ”francesco” ”mosto” collects from his entire real estate portfolio?
53 What types of properties can be found in ”calle di santa cattarina principia alle fondamenta nove”?
@@ 56 Which properties generate a rental income exceeding 50 ducati?
57 In each location, what is the count of ”casa” properties?
@@ 65 Who are the individuals possessing properties in several areas?
66 Is there availability of ”casa” type properties located on ”calle della malvasia”?
67 Who are the proprietors of a ”bottega” located at ”fondamenta in fazza la beccaria”?
68 How many properties out of a hundred are found in ”fondamenta de carmini”?
69 What’s the minimum amount ”zuane” ”panizza” earns from the property at ”traghetto di san silvestro”?
70 What’s the typical rental income for each type of property owned by ”perina” ”capello”?
71 What is the count of properties classified under ”casa a pepian”, and how do they contribute to the overall rental revenue percentage-wise in the data?
@@ 72 In the dataset, which entity holds ownership over the largest number of properties?
73 Does ”antonio” possess any real estate on ”calle dei ragusei”?
@@ 75 How is the dataset broken down by property type?
76 What is the count of proprietors possessing a minimum of one ”porzione di bottega” property?
@@ 78 Who is the principal owner of the most varied property categories?
79 Is a property in ”campiello della fraterna” owned by ”iseppo maria” ”gallo”?
80 Who holds the title as the largest landholder in ”sotto le collonelle”?
81 How many properties do owners have on average in the location of ”corte carli”?
@@ 82 What’s the total of property owners with investments spread across several areas?
83 In ”loco vicino la calle del paradiso”, are there identical property types under the ownership of different families?
84 Is the property yielding the least rental income in the dataset owned by ”domenico” ”michiel”?
85 How does the total earnings from rent of the ”casetta, vuota, di solito affittata” stand in comparison with that from different kinds of properties?
@@ 87 How much total rent income do properties with more than 10 ducati generate?
88 How many families own property in ”calle corrente dell’occa”?
89 What is the lowest income of ”francesco” ”giustinian”?
90 For rental properties, what share of the income is derived from those labeled ”bottega da fabro”?
@@ 96 How many owners have more than one type of property?
97 How is rent income distributed among properties in ”rio terrà”?
@@ A.2 SQL CodeS-7B Prompts
A.3 Coding Agent Implementation Details
@@
A.4 Coding Agent Prompts
This section contains all the prompts used at each step of the coding agent’s process. The first two prompts serve as system prompts, provided to the agent along with a description of the datasets as part of its context. The subsequent prompts align with the agent’s workflow, as shown in Figure 5. The process begins with extracting references, followed by entity extraction. Next, the agent generates a plan and subsequently writes the necessary code. If errors occur during code execution, the agent debugs the code based on the Python console’s error messages. In the following prompts, all inputs used to generate prompts are highlighted in blue. In context examples are given for reference and entity extraction.